OpenStudy (anonymous):

The vertices of a triangle are A(16,0) B(9,2) and C(0,0) which side is also an altitude with a slope of -2/7?

OpenStudy (praxer):

it is BA. $$\huge slope = \dfrac{y_2-y_1} {x_2-x_1}$$ plug in the values for BA you will get, $$\huge slope =\dfrac{0-2}{16-9}$$ $$\huge it\ is = -2/7$$

OpenStudy (anonymous):

thank you, geometry is so hard for me

OpenStudy (praxer):

$$\huge \frak\color{red} {WELCOME..........}$$